OpenStudy (iwanttogotostanford):

Really Need Help Here, Clueless: Describe how the graph of y= x^2 can be transformed to the graph of the given equation. y = x^2 - 14

OpenStudy (iwanttogotostanford):

@phi

OpenStudy (agent0smith):

You shouldn't be clueless; you must have some notes related to this or something. Eg. y = x^2 + 3 can be obtained by shifting the graph of y = x^2 up 3 units.

OpenStudy (iwanttogotostanford):

well i am, I started learning it but I still don't get it. So I am thinking it would be: Shift the graph of y = x^2 down 14 units ??

OpenStudy (iwanttogotostanford):

?

OpenStudy (agent0smith):

Yes

OpenStudy (iwanttogotostanford):

and I am confused on one more. It is: Describe how the graph of y= x^2 can be transformed to the graph of the given equation. y = (x-14)^2 - 9. I am confused to if it would be: A. Shift the graph of y = x^2 down 14 units and then left 9 units or B. Shift the graph of y = x^2 left 14 units and then down 9 units. how would I figure out which one? I can't decide

OpenStudy (agent0smith):

What if it just said y= x^2 - 9?

OpenStudy (agent0smith):

Neither of those are right btw

OpenStudy (phi):

if you start with y= x^2 you "put in an x" and "get out a y" if you added -14 to the equation: y= x^2 -14 the "new y" is less than the old y (by 14) if you plotted the "old y" and the new y, the new y is 14 steps below the old y

OpenStudy (phi):

for your problem, where the new y is y = (x-14)^2 - 9. which shows -9 being added. that shifts the graph which way compared to y=(x-14)^2 ?

OpenStudy (phi):

when you "add -14" to the x (i.e. change x^2 to (x-14)^2 ) that shifts the graph sideways (by 14). The only problem is it's "backwards" (so very confusing) x-14 shifts the graph 14 to the right (and x+14 would shift 14 to the left)

OpenStudy (iwanttogotostanford):

@phi @agent0smith oh wow and I forgot it was a horizontal translation. I figured out that it would shift to the right actually and not left since it is a horizontal translation. SO, my answer is (what I think, anyways): Shift the graph of y = x^2 right 14 units and then down 9 units.

OpenStudy (iwanttogotostanford):

@agent0smith

OpenStudy (iwanttogotostanford):

@agent0smith hello?

OpenStudy (iwanttogotostanford):

@agent0smith

OpenStudy (iwanttogotostanford):

@mathstudent55 @imqwerty